which of the following shows the correct procedure to find the Present value of $100,000 received at the end of 5 years, if the discount rate is 4%?
Answer #1
Present value of cash flow = Future value of cash flow * Present value of 1
= $       1,00,000 * 0.821927
= $     82,192.71
Working:
Present value of 1 = (1+i)^-n Where,
= (1+0.04)^-5 i = 4%
= 0.821927107 n = 5